# CGKN: A Deep Learning Framework for Modeling Complex Dynamical Systems and Efficient Data Assimilation

### v1

@@ 3.2 Lorenz 96 system: A forty-dimensional chaotic system
In this test case, the parameter $J$ is selected as 6, and therefore the dimension of latent state $d_{\mathbf{v}}$ in Eq.(2.3) is $12020J=$ 120. The $\textrm{NN}_{i}^{(\mathbf{u}_{1})}$ stands for the $(i+1)$-th output of the network $\textrm{NN}^{(\mathbf{u}_{1})}\in\mathbb{R}^{3}\mapsto\mathbb{R}^{19}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}x_{2i-1}/\mathrm{d}t$ and similar to $\textrm{NN}^{(\mathbf{v}_{j})}\in\mathbb{R}^{3}\mapsto\mathbb{R}^{31}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}\mathbf{v}^{(i)}_{j}/\mathrm{d}t$. Since all the neural networks $\textrm{NN}^{(\mathbf{v}_{j})}$ take the same inputs for each $j=1,2,\cdots,J$, they are merged into a single neural network $\textrm{NN}^{(\mathbf{v})}:\in\mathbb{R}^{3}\mapsto\mathbb{R}^{186}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}\mathbf{v}^{(i)}/\mathrm{d}t=[\mathrm{d}\mathbf{v}_{1}^{(i)}/\mathrm{d}t,\mathrm{d}\mathbf{v}_{2}^{(i)}/\mathrm{d}t,\cdots,\mathrm{d}\mathbf{v}_{J}^{(i)}/\mathrm{d}t]$. When iterating $\textrm{NN}^{(\mathbf{u}_{1})}$ across all dynamics of observed state, the outputs will formalize the $\mathbf{f}_{1}\in\mathbb{R}^{20\times 1}$ and $\mathbf{g}_{1}\in\mathbb{R}^{20\times 120}$ in (2.3). Similarly, iterating $\textrm{NN}^{(\mathbf{v})}$ across all dynamics of latent state will formalize the $\mathbf{f}_{2}\in\mathbb{R}^{120\times 1}$ and $\mathbf{g}_{2}\in\mathbb{R}^{120\times 120}$. One thing should be noted is that, due to the locality strategy, the $\mathbf{f}_{2}$ and $\mathbf{g}_{2}$ are sparse matrices with zero elements representing non-interaction with the uncorrelated states and non-zero elements filled by the outputs of $\textrm{NN}^{(\mathbf{u}_{1})}$ and $\textrm{NN}^{(\mathbf{v})}$ respectively. The sub-networks in the CGKN with locality and homogeneity of this test case are $\bm{\varphi}:\mathbb{R}^{2}\mapsto\mathbb{R}^{6}$, $\bm{\psi}:\mathbb{R}^{6}\mapsto\mathbb{R}^{2}$ and $\bm{\eta}=[\textrm{NN}^{(\mathbf{u}_{1})},\textrm{NN}^{(\mathbf{v})}]$ with the number of parameters $\bm{\theta}_{\bm{\varphi}}$, $\bm{\theta}_{\bm{\psi}}$ and $\bm{\theta}_{\bm{\eta}}$ be 8646, 8942 and 5757, respectively. The training setup for CGKN is as follows: the state forecast horizon is $t_{N_{s}}=0.2$ time units, the DA horizon is $t_{N_{l}}=60$ time units with the first $t_{N_{b}}=2$ time units be cut out. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=\lambda_{\mathbf{v}}=1/d_{\mathbf{u}}$, and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is selected as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler. It is worth noting that the complexity of the CGKN incorporated with the locality and homogeneity does not increase with the dimensionality of the Lorenz 96 system which makes CGKN scalable to even higher dimensions.
@@
The test results for the different models are reported in Table 3.1. For the true model, we use the mean of ensemble simulations for state prediction and employ the EnKBF for DA, with an ensemble size of 100 for both tasks. Among all the other models, the CGKN is the only one that achieves similar performance in both state forecasting and DA compared to the true model. In terms of state forecasting, the DNN model performs slightly better than the CGKN model. However, the CGKN’s advantage lies in its ability to avoid the computationally expensive DA techniques typically required for nonlinear models, making it significantly more efficient for DA. Regarding models that allow efficient DA, the simplified KoopNet model slightly outperforms the CG-Reg model; however, both models significantly underperform compared to the CGKN model.
@@
Figure 3.4 illustrates the true state variables alongside the prediction results from different models with a lead time of 0.2 time units. The Hovmoller diagram (namely, the spatiotemporal field) of the true system and the simulation results from the CGKN model indicate that the CGKN successfully captures the overall chaotic behavior and wave patterns. The time series of state variables $x_{1}$ and $x_{2}$ from the true system, along with the forecast results from the various models, are also presented in Fig. 3.4. The results for other observable and unobserved state variables are similar to those of $x_{1}$ and $x_{2}$ due to the statistical homogeneity of this test case and are thus omitted. Notably, both the simplified KoopNet model and the CG-Reg model exhibit significant discrepancies compared to the truth, while the CGKN model shows results that are more accurate. Although the simplified KoopNet model performs reasonably well in forecasting the observable variable $x_{1}$, its performance for the unobserved state variable $x_{2}$ is inferior to that of the CGKN model. This discrepancy primarily arises from the neglect of the contribution of observed state variables to the dynamics of unobserved state variables in the simplified KoopNet model, which is a crucial factor in the true system due to the quadratic term in Eq. (3.8).
@@
Figure 3.4: State forecast results with a lead time of 0.2 time units. Panel (a) and (b) are the Hovmoller diagram of true system and results from the CGKN model. The bottom two panels show the true signal of states $x_{1}$ and $x_{2}$ and the corresponding results from three different models.
@@
Figure 3.5 presents the DA results of the unobserved state $x_{2}$. In each panel, the lower sub-figure provides a zoomed-in view of the upper one to clearly display the trajectory and uncertain area. In this case, the DA results from the true model with EnKBF yield a posterior mean that closely aligns with the true signal and a narrow uncertainty area, which is obtained from the estimated covariance of ensemble simulations. The CGKN model shows a good agreement with the true model and outperforms the simplified KoopNet model and the CG-Reg model. For the CG-Reg model, the uncertainty area is derived from the covariance evolution formula with a similar form of Eq. (LABEL:eq:CGKN_CGF). For both the CGKN model and the simplified KoopNet model, the uncertainty area is based on the residual analysis method described in Section 2.1.2.
@@
Figure 3.5: DA results of the unobserved state variable $x_{2}$ in the Lorenz 96 system. EnKBF is used for the true model, and analytic DA formulae are used for the other three models. The uncertainties are indicated by the grey-colored regions, which correspond to two estimated standard deviations from the posterior mean. The uncertainty area from the CGKN model and the simplified KoopNet model are derived from residual analysis.
@@ 3.3 El Niño-Southern Oscillation (ENSO): A multiscale non-Gaussian phenomenon
A physics-based reference model is employed to generate synthetic ENSO data [110] for training and testing different methods here. This model consists of a set of stochastic partial differential equations (SPDEs) that describe the coupled dynamics of the atmosphere and ocean. It effectively captures many realistic features observed in the real world, including non-Gaussian statistics and variations in spatial patterns, peak intensity, and temporal evolution. Its capability to generate much longer time series than actual observations enables us to train and evaluate the skill of the CGKN on this complex natural phenomenon.
@@
Figure 3.6 displays a portion of the training data for $T_{E}$, $H_{W}$, and $a_{p}$ in panel (a), along with their corresponding PDFs and ACFs in panels (b) and (c), evaluated using the entire training dataset. In this ENSO test case, the observed state variables are $T_{W}$, $T_{C}$, $T_{E}$, and $a_{p}$, while the unobservable ones are $H_{W}$, $H_{C}$, $H_{E}$, with $a_{p}$ serving as an external variable. The models, including CGKN, are trained on the data of all state variables and aim to forecast both $T$ and $H$ states, while performing DA without access to $H$ data. In practice, SST data is obtained from satellite instruments, while directly measuring thermocline depth is challenging. DA is involved in recovering the thermocline depth in practice. Therefore, estimating thermocline depth from SST in this setup simulates a realistic scenario.
@@
where $\mathbf{v}=\bm{\varphi}(H_{W},H_{C},H_{E})$ and $[H_{W},H_{C},H_{E}]^{\mathtt{T}}=\bm{\psi}(\mathbf{v})$. The latent state dimension of the autoencoder is set as 10 in this test case. With this choice, the sub-networks in the CGKN model are $\bm{\varphi}:\mathbb{R}^{3}\mapsto\mathbb{R}^{10}$, $\bm{\psi}:\mathbb{R}^{10}\mapsto\mathbb{R}^{3}$, $\bm{\eta}:\mathbb{R}^{4}\mapsto\mathbb{R}^{154}$ with number of parameters being 10554, 12970, 12963, respectively. For the training settings, the state forecast time range $[t_{0},t_{N_{s}}]$ is set as 1 year, and DA time range $[t_{0},t_{N_{l}}]$ is 40 years with the warm-up period $t_{N_{b}}$ being 4 years. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=1/d_{\mathbf{u}}$, $\lambda_{\mathbf{v}}=0$ and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is selected as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler.
@@
The test results of all models for this case has been reported in the Table 3.1. The ensemble number of the true model for ensemble mean state forecast is set to 100. The DA error from EnKBF has been omitted due to the computational resource required by the ENSO PDEs and it is not the focus of this work. For the performance of state forecast, except for the CG-Reg model, the other models achieve similar results with the simplified KoopNet model ranked the least. For the DA performance, the CG-Reg model becomes unstable and diverges over the horizon of 2000 years from test data. The DA error of the CGKN model is the lowest, while that of the simplified KoopNet model is approximately twice as large. In this test case, the CGKN model demonstrates optimal performance in both state forecast and DA compared to all other models.
@@
Figure 3.7 shows the lead time prediction of different models in terms of $T_{E}$, which is one of the most widely used ENSO indicator. The panel (a) and (b) display the NRMSE and the pattern correlation between the true signal of state $T_{E}$ and its predictive states at different lead times. The correlation is a standard measurement in climate science and it is therefore included here. The NRMSE is calculated based on the formula of Eq. (3.4) and the correlation is from the formula:
$\textrm{Corr}=\frac{\sum_{n=1}^{N}\big{(}x^{\dagger}(t_{n})-\bar{x}^{\dagger}\big{)}\big{(}x(t_{n})-\bar{x}\big{)}}{\sqrt{\sum_{n=1}^{N}\big{(}x^{\dagger}(t_{n})-\bar{x}^{\dagger}\big{)}^{2}}\sqrt{\sum_{n=1}^{N}\big{(}x(t_{n})-\bar{x}\big{)}^{2}}}$ (3.11)
@@
with $x^{\dagger}\in\mathbb{R}^{N}$ is the true values of state time series, $x$ is the approximated values, $x(t_{n})$ is the state at time $t_{n}$. The $\bar{x}^{\dagger}$ and $\bar{x}$ are the time average of true values and approximated values. The panels showing NRMSE and correlation indicate that both the CGKN model and the simplified KoopNet model produce stable and accurate state forecast results across various lead times, demonstrating the effectiveness of these deep learning models. In contrast, the CG-Reg model exhibits a rapid decline in forecast performance due to its omission of certain key dynamics in the parametric forms, while the persistence model performs the worst among all the models. Panels (c)-(d) show the time series of the true state and predictive states at a lead time of 3-month, 6-month, 9-month, and 12-month, respectively. For the lead time prediction at 3-month in panel (c), all the models including persistence can almost follow the true signal. However, with the lead time becoming longer, the deviation and shift pattern of predictions from all models have become more pronounced. For the 12-month lead time prediction in panel (f), although the forecast skill of the CGKN model declines, it still performs the best overall. The model maintains a correlation above 0.5 and significantly outperforms the persistence model.
@@
Figure 3.7: Comparison of the lead-time forecast for state $T_{E}$ from different models. Panel (a) and (b): skill scores including normalized root mean squared error (NRMSE) and correlation of the CGKN model, simplified KoopNet model, CG-Reg model, and Persistence. Panel (c) - (f): comparison of the true signal and forecast results from different models at lead times of 3, 6, 9, and 12 months, respectively, over the years 1020 to 1050.
@@
Figure 3.8 shows the DA results from the CGKN model for estimating the unobserved state $H_{W}$, $H_{C}$ and $H_{E}$ given trajectories of observed state $T_{W}$, $T_{C}$, $T_{E}$ and $a_{p}$. The DA results of the true model have been omitted due to the complexity of the governing SPDEs and because the DA results from EnKBF are not the main focus of this work. The DA posterior mean from the CGKN model is obtained by applying the analytic formulae in Eq. (LABEL:eq:CGKN_CGF), while the uncertainty areas are determined using the method of residual analysis described in Section 2.1.2. The similarity between the posterior mean from the CGKN model and the true signal confirms its capability in DA. Although the posterior mean from the CGKN model shows does not perfectly match the true signal occasionally, the uncertainty area can still cover the true data for most of those times.
@@
Figure 3.9: Hovmoller diagrams for the simulation from true model and reconstructions by true and approximated states. Panel (a): simulation of $\mathbf{T}$ (SST) from the true model. Panel (b): reconstruction of $\mathbf{T}$ from the true discrete states $T_{W}$, $T_{C}$ and $T_{E}$ which are spatial average of $\mathbf{T}$. Panel (c): reconstruction of $\mathbf{T}$ from predictive states $T_{W}$, $T_{C}$ and $T_{E}$ of the CGKN model at 6-month lead time. Panel (d): simulation of $\mathbf{H}$ (thermocline) from the true model. Panel (e): reconstruction of $\mathbf{H}$ from the true discrete states $H_{W}$, $H_{C}$ and $H_{E}$ which are spatial average of $\mathbf{H}$. Panel (f): reconstruction of $\mathbf{H}$ from estimated $H_{W}$, $H_{C}$, and $H_{E}$ via DA which apply EnKBF on CGKN.

### v2

@@ 3.2 Lorenz 96 system: A forty-dimensional chaotic system
In this test case, the parameter $J$ is selected as 6, and therefore, the dimension of latent state $d_{\mathbf{v}}$ in Eq.(2.8) is $12020J=$ 120. The $\textrm{NN}_{i}^{(\mathbf{u}_{1})}$ stands for the $(i+1)$-th output of the network $\textrm{NN}^{(\mathbf{u}_{1})}\in\mathbb{R}^{3}\mapsto\mathbb{R}^{19}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}x_{2i-1}/\mathrm{d}t$ and similar to $\textrm{NN}^{(\mathbf{v}_{j})}\in\mathbb{R}^{3}\mapsto\mathbb{R}^{31}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}\mathbf{v}^{(i)}_{j}/\mathrm{d}t$. Since all the neural networks $\textrm{NN}^{(\mathbf{v}_{j})}$ take the same inputs for each $j=1,2,\cdots,J$, they are merged into a single neural network $\textrm{NN}^{(\mathbf{v})}:\in\mathbb{R}^{3}\mapsto\mathbb{R}^{186}$ which takes input $(x_{2i-3},x_{2i-1},x_{2i+1})$ for dynamics $\mathrm{d}\mathbf{v}^{(i)}/\mathrm{d}t=[\mathrm{d}\mathbf{v}_{1}^{(i)}/\mathrm{d}t,\mathrm{d}\mathbf{v}_{2}^{(i)}/\mathrm{d}t,\cdots,\mathrm{d}\mathbf{v}_{J}^{(i)}/\mathrm{d}t]$. When iterating $\textrm{NN}^{(\mathbf{u}_{1})}$ across all dynamics of observed state, the outputs will formalize the $\mathbf{f}_{1}\in\mathbb{R}^{20\times 1}$ and $\mathbf{g}_{1}\in\mathbb{R}^{20\times 120}$ in Eq. (2.8). Similarly, iterating $\textrm{NN}^{(\mathbf{v})}$ across all dynamics of latent state will formalize the $\mathbf{f}_{2}\in\mathbb{R}^{120\times 1}$ and $\mathbf{g}_{2}\in\mathbb{R}^{120\times 120}$. One thing should be noted is that, due to the locality strategy, the $\mathbf{f}_{2}$ and $\mathbf{g}_{2}$ are sparse matrices with zero elements representing non-interaction with the uncorrelated states and non-zero elements filled by the outputs of $\textrm{NN}^{(\mathbf{u}_{1})}$ and $\textrm{NN}^{(\mathbf{v})}$ respectively. The sub-networks in the CGKN with locality and homogeneity of this test case are $\bm{\varphi}:\mathbb{R}^{2}\mapsto\mathbb{R}^{6}$, $\bm{\psi}:\mathbb{R}^{6}\mapsto\mathbb{R}^{2}$ and $\bm{\eta}=[\textrm{NN}^{(\mathbf{u}_{1})},\textrm{NN}^{(\mathbf{v})}]$ with the number of parameters $\bm{\theta}_{\bm{\varphi}}$, $\bm{\theta}_{\bm{\psi}}$ and $\bm{\theta}_{\bm{\eta}}$ be 8646, 8942 and 5757, respectively. In the CGKN, the encoder $\bm{\varphi}$, decoder $\bm{\psi}$ and sub-networks $\bm{\eta}$ including $\textrm{NN}^{(\mathbf{v})}$ and $\textrm{NN}^{(\mathbf{u}_{1})}$ are all fully connected networks with 5, 5, 4, and 4 layers, respectively. The activation function is chosen as ReLU. The training setup for CGKN is as follows: the state forecast horizon is $t_{N_{s}}=0.2$ time units, the DA horizon is $t_{N_{l}}=60$ time units with the first $t_{N_{b}}=2$ time units be cut out. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=\lambda_{\mathbf{v}}=1/d_{\mathbf{u}}$, and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is selected as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler. It is worth noting that the complexity of the CGKN incorporated with the locality and homogeneity does not increase with the dimensionality of the Lorenz 96 system which makes CGKN scalable to even higher dimensions.
Table 3.2: Test results of state forecast and data assimilation of all models for the example of Lorenz 96 system. The errors are the normalized root mean squared error (NRMSE) between true values and approximated values.
Models Errors Forecast Error DA Error True Model 4.3874e-02 6.9582e-02 CGKN 6.3738e-02 9.4883e-02 Standard KoopNet 3.0589e-01 2.7252e-0 CG-Reg 2.7631e-01 3.5670e-01 DNN 4.5174e-02
@@
The test results for the different models are reported in Table 3.2. For the true model, we use the mean of ensemble simulations for state prediction and employ the EnKBF for DA, with an ensemble size of 100 for both tasks. Among all the other models, the CGKN is the only one that achieves similar performance in both state forecast and DA compared to the true model. In terms of state forecast, the DNN model performs slightly better than the CGKN model. However, the CGKN’s advantage lies in its ability to avoid the computationally expensive DA techniques typically required for nonlinear models, making it significantly more efficient for DA. Regarding models that allow efficient DA, the standard KoopNet model slightly outperforms the CG-Reg model; however, both models underperform compared to the CGKN model.
@@
Figure 3.4 illustrates the true state variables alongside the prediction results from different models with a lead time of 0.2 time units. The Hovmoller diagram (namely, the spatiotemporal field) of the true system and the simulation results from the CGKN model indicate that the CGKN successfully captures the overall chaotic behavior and wave patterns. The time series of state variables $x_{1}$ and $x_{2}$ from the true system, along with the forecast results from the various models, are also presented in Fig. 3.4. The results for other observed and unobserved state variables are similar to those of $x_{1}$ and $x_{2}$ due to the statistical homogeneity of this test case and are thus omitted. Notably, both the standard KoopNet model and the CG-Reg model exhibit discrepancies compared to the truth, while the CGKN model shows results that are more accurate. Although the standard KoopNet model performs reasonably well in forecasting the observed variable $x_{1}$, its performance for the unobserved state variable $x_{2}$ is inferior to that of the CGKN model. This discrepancy primarily arises from the neglect of the contribution of observed state variables to the dynamics of unobserved state variables in the standard KoopNet model, which is a crucial factor in the true system due to the quadratic term in Eq. (3.8).
@@
Figure 3.4: State forecast results with a lead time of 0.2 time units. Panel (a) and (b) are the Hovmoller diagram of the true system and results from the CGKN model. The bottom two panels show the true signal of states $x_{1}$ and $x_{2}$ and the corresponding results from three different models.
@@
Figure 3.5 presents the DA results of the unobserved state $x_{2}$. In each panel, the lower sub-figure provides a zoomed-in view of the upper one to clearly display the trajectory and uncertain area. In this case, the DA results from the true model with EnKBF yield a posterior mean that closely aligns with the true signal and a narrow uncertainty area, which is obtained from the estimated covariance of ensemble simulations. The CGKN model shows a good agreement with the true model and outperforms the standard KoopNet model and the CG-Reg model. For the CG-Reg model, the uncertainty area is derived from the covariance evolution formula with a similar form of Eq. (LABEL:eq:CGKN_CGF). For both the CGKN model and the standard KoopNet model, the uncertainty area is based on the residual analysis method described in Section 2.2.2.
@@
Figure 3.5: DA results of the unobserved state variable $x_{2}$ in the Lorenz 96 system. EnKBF is used for the true model, and analytic DA formulae are used for the other three models. The uncertainties are indicated by the grey-colored regions, which correspond to two estimated standard deviations from the posterior mean. The uncertainty area from the CGKN model and the standard KoopNet model are derived from residual analysis.
@@ 3.3 El Niño-Southern Oscillation (ENSO): A multiscale non-Gaussian phenomenon
A physics-based reference model is employed to generate synthetic ENSO data [113] for training and testing different methods here. This model consists of a set of stochastic partial differential equations (SPDEs) that describe the coupled dynamics of the atmosphere and ocean. It effectively captures many realistic features observed in the real world, including non-Gaussian statistics and variations in spatial patterns, peak intensity, and temporal evolution. Its capability to generate much longer time series than actual observations enables us to train and evaluate the skill of the CGKN on this complex natural phenomenon. More details and the governing equations of the ENSO SPDEs are available in B.
@@
Figure 3.6 displays a portion of the training data for $T_{E}$, $H_{W}$, and $a_{p}$ in panel (a), along with their corresponding PDFs and ACFs in panels (b) and (c), evaluated using the entire training dataset. In this ENSO test case, the observed state variables are $T_{W}$, $T_{C}$, $T_{E}$, and $a_{p}$, while the unobserved ones are $H_{W}$, $H_{C}$, $H_{E}$, with $a_{p}$ serving as an external variable. The models, including CGKN, are trained on the data of all state variables and aim to forecast both $T$ and $H$ states, while performing DA without access to $H$ data. In practice, SST data is obtained from satellite instruments, while directly measuring thermocline depth is challenging. DA is involved in recovering the thermocline depth in practice. Therefore, estimating thermocline depth from SST in this setup simulates a realistic scenario.
@@
where $\mathbf{v}=\bm{\varphi}(H_{W},H_{C},H_{E})$ and $[H_{W},H_{C},H_{E}]^{\mathtt{T}}=\bm{\psi}(\mathbf{v})$. The latent state dimension of the autoencoder is set as 10 in this test case. With this choice, the sub-networks in the CGKN model are $\bm{\varphi}:\mathbb{R}^{3}\mapsto\mathbb{R}^{10}$, $\bm{\psi}:\mathbb{R}^{10}\mapsto\mathbb{R}^{3}$, $\bm{\eta}:\mathbb{R}^{4}\mapsto\mathbb{R}^{154}$ with number of parameters being 10554, 12970, 12963, respectively. In CGKN, the fully connected neural networks are used for encoder $\bm{\varphi}$, decoder $\bm{\psi}$, and sub-network $\bm{\eta}$, each consisting of six layers with ReLU activation functions. For the training settings, the state forecast time range $[t_{0},t_{N_{s}}]$ is set as 1 year, and DA time range $[t_{0},t_{N_{l}}]$ is 40 years with the warm-up period $t_{N_{b}}$ being 4 years. The weights in the target loss function are set as $\lambda_{\textrm{AE}}=1/d_{\mathbf{u}_{2}}$, $\lambda_{\mathbf{u}}=1/d_{\mathbf{u}}$, $\lambda_{\mathbf{v}}=0$ and $\lambda_{\textrm{DA}}=1/d_{\mathbf{u}_{2}}$. The gradient descent optimizer is selected as Adam with an initial learning rate of 1e-3 and a cosine annealing scheduler.
Table 3.3: Test results of state forecast and data assimilation of all models for the example of ENSO. The errors are the normalized root mean squared error (NRMSE) between true values and approximated values.
Models Errors Forecast Error DA Error True Model 6.5406e-01 CGKN 6.4762e-01 1.3518e-01 Standard KoopNet 6.8473e-01 2.3490e-01 CG-Reg 9.8751e-01 NaN DNN 6.4845e-01
@@
The test results of all models for this case have been reported in Table 3.3. The ensemble number of the true model for ensemble mean state forecast is set to 100. The DA error from EnKBF has been omitted due to the computational resource required by the ENSO PDEs and it is not the focus of this work. For the performance of state forecast, except for the CG-Reg model, the other models achieve similar results with the standard KoopNet model ranked the lowest. For the DA performance, the CG-Reg model becomes unstable and diverges over the horizon of 2000 years from test data. The DA error of the CGKN model is the lowest, outperforming the performance of the standard KoopNet. In this test case, the CGKN model demonstrates optimal performance in both state forecast and DA compared to all other models.
@@
Figure 3.7 shows the lead time prediction of different models in terms of $T_{E}$, which is one of the most widely used ENSO indicator. The panels (a) and (b) display the NRMSE and the pattern correlation between the true signal of state $T_{E}$ and its predictive states at different lead times. The correlation is a standard measurement in climate science and it is therefore included here. The NRMSE is calculated based on the formula of Eq. (3.4) and the correlation is from the formula:
$\textrm{Corr}=\frac{\sum_{n=1}^{N}\big{(}x^{\star}(t_{n})-\bar{x}^{\star}\big{)}\big{(}x(t_{n})-\bar{x}\big{)}}{\sqrt{\sum_{n=1}^{N}\big{(}x^{\star}(t_{n})-\bar{x}^{\star}\big{)}^{2}}\sqrt{\sum_{n=1}^{N}\big{(}x(t_{n})-\bar{x}\big{)}^{2}}}$ (3.11)
@@
with $x^{\star}\in\mathbb{R}^{N}$ is the true values of state time series, $x$ is the approximated values, $x(t_{n})$ is the state at time $t_{n}$. The $\bar{x}^{\star}$ and $\bar{x}$ are the time average of true values and approximated values. In this example, the persistence model, a simple and commonly used prediction method for ENSO, is considered. It assumes that ENSO conditions remain unchanged over time and use the current values as future predictions. The persistence serves as a baseline model for evaluating other computational models. The panels showing NRMSE and correlation indicate that both the CGKN model and the standard KoopNet model produce stable and accurate state forecast results across various lead times, demonstrating the effectiveness of these deep learning models. In contrast, the CG-Reg model exhibits a rapid decline in forecast performance due to its omission of certain key dynamics in the parametric forms, while the persistence model performs the worst among all the models. Panels (c)-(d) show the time series of the true state and predictive states at a lead time of 3 months, 6 months, 9 months, and 12 months, respectively. For the lead time prediction at 3 months in panel (c), all the models including persistence can almost follow the true signal. However, with the lead time becoming longer, the deviation and shift pattern of predictions from all models have become more pronounced. For the 12-month lead time prediction in panel (f), although the forecast skill of the CGKN model declines, it still performs the best overall. The model maintains a correlation above 0.5 and significantly outperforms the persistence model.
@@
Figure 3.7: Comparison of the lead-time forecast for state $T_{E}$ from different models. Panel (a) and (b): skill scores including normalized root mean squared error (NRMSE) and correlation of the CGKN model, standard KoopNet model, CG-Reg model, and Persistence. Panel (c) - (f): comparison of the true signal and forecast results from different models at lead times of 3, 6, 9, and 12 months, respectively, over the years 1020 to 1050.
@@
Figure 3.8 shows the DA results from the CGKN model for estimating the unobserved state $H_{W}$, $H_{C}$ and $H_{E}$ given trajectories of observed state $T_{W}$, $T_{C}$, $T_{E}$ and $a_{p}$. The governing equations of the true model for ENSO are stochastic partial differential equations (SPDEs), detailed in Appendix B. The SPDEs are computationally intensive and time-consuming for ensemble methods to perform data assimilation effectively. For the CG-Reg model, the DA result is very unstable and diverges quickly, leading to the finite-time blow-up of the posterior mean and covariance. In addition, the standard KoopNet shows a similar DA result to the CGKN. Therefore, Figure 3.8 primarily focuses on the DA results of CGKN compared to the true signal. The DA posterior mean from the CGKN model is obtained by applying the analytic formulae in Eq. (LABEL:eq:CGKN_CGF), while the uncertainty areas are determined using the method of residual analysis described in Section 2.2.2. The similarity between the posterior mean from the CGKN model and the true signal confirms its capability in DA. Although the posterior mean from the CGKN model shows does not perfectly match the true signal occasionally, the uncertainty area can still cover the true data for most of those times.
@@
Figure 3.9: Hovmoller diagrams for the simulation from true model and reconstructions by true and approximated states. Panel (a): simulation of $\mathbf{T}$ (SST) from the true model. Panel (b): reconstruction of $\mathbf{T}$ from the true discrete states $T_{W}$, $T_{C}$ and $T_{E}$ which are spatial average of $\mathbf{T}$. Panel (c): reconstruction of $\mathbf{T}$ from predictive states $T_{W}$, $T_{C}$ and $T_{E}$ of the CGKN model at a 6-month lead time. Panel (d): simulation of $\mathbf{H}$ (thermocline) from the true model. Panel (e): reconstruction of $\mathbf{H}$ from the true discrete states $H_{W}$, $H_{C}$ and $H_{E}$ which are spatial average of $\mathbf{H}$. Panel (f): reconstruction of $\mathbf{H}$ from estimated $H_{W}$, $H_{C}$, and $H_{E}$ via DA which apply analytical formulae on CGKN.